One- Deminsion Kinematics

1. The marathon runner Jogged 2 km along a straight shoreline, rested, then continued in the same direction for another 2 km. Find the total distance he covered and his total
displacement

2. Timothy drives 10 km to the South and then 20 km to the East. What is the magnitude and direction of his displacement from the origin?

3. What is the difference between distance and displacement?

4. How can velocity be obtained?

5. Can the displacement of an object affect its velocity?

6. Can the direction have a big impact on the value of velocity?

7 Is acceleration dependent on velocity? How?

8. Can a change in velocity affect the value of acceleration?

9. What is the relationship between acceleration and velocity?
1
Expert's answer
2021-09-28T11:14:18-0400

1)


"s=d=4\ km"

2)


"s=\sqrt{10^2+20^2}=22.4\ km\\\theta=\arctan{\frac{10}{20}}=26.6\degree\text{ S of E}"

3) Distance is a scalar quantity that refers to "how much ground an object has covered" during its motion. Displacement is a vector quantity that refers to "how far out of place an object is"; it is the object's overall change in position.

4) We can obtain an expression for velocity as the area under an a(t) acceleration vs. time graph.

5) No.

6) Yes since velocity is a vector, it involves magnitude and direction.

7) No

8)  The acceleration of the object is dependent upon this velocity change and is in the same direction as this velocity change. The acceleration of the object is in the same direction as the velocity change vector.

9) Velocity is the rate of change of displacement. Acceleration is the rate of change of velocity. Velocity is a vector quantity because it consists of both magnitude and direction. Acceleration is also a vector quantity as it is just the rate of change of velocity.